HUBBLE SPACE TELESCOPE DAILY REPORT #1761 PERIOD COVERED: 0000Z (UTC) 11/15/96 - 0000Z (UTC) 11/18/96 Daily Status Report as of 323/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive = Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.2 Completed Four Sets of FOS/BL 6407 (Post Asymptotic Giant Branch = Evolution in the Magellanic Clouds) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet spectrographic observations of planetary nebulae LMC-SMP15, = LMC-SMP41, LMC-SMP54 and LMC-SMP01 which are located in the Large = Magellanic Cloud. These data should give the spatial structure, size, = ionized masses, and dynamic ages of these objects. At 321/052249Z, = during the second iteration of this proposal, there was an FOS target = acquisition failure. HSTAR 5799 was written. The observations were = executed as planned, and no problems were reported. 1.3 Completed FOS/BL 6683 (Eclipsing Binaries in the Magellanic Clouds = - Fundamental Properties A) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of O/B Class, eclipsing binary system star = LMC HV2274 located in the Large Magellanic Cloud. From these data and = the latest Kurucz model atmospheres, accurate temperatures will be = derived. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed FOS/RD 6654 (Ultraviolet Spectroscopy of Hot DB White = Dwarfs) The Faint Object Spectrograph (Red Side) was used to make = observations of the hot, helium-atmosphere, DB White Dwarf class, star = WD0112+104. The ultraviolet spectra will be used to determine the = effective temperatures of this star, which shows discrepancies when = calculated from the optical spectra due to traces of photospheric = hydrogen which is not visible at optical wavelengths. The observations = were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 7211 (WF/PC-2 Pre-SM97 UV Baseline = Observation.) The WF/PC-2 was used to make far-ultraviolet observations of CVZ = Class targets GRW+70D582 and WD0310-688 as part of a pre-observation = baseline of SM97 (WF/PC-2 SMOV Calibration). The observations were = completed as scheduled, and no anomalies were noted. 1.6 Completed FOS/BL 6546 (Dissecting the wind and Disk of the = Nova-like Variable V347 Puppus) The Faint Object Spectrograph (Blue Side) was used to make = observations of the nova-like, variable star V347 Puppus in order to = study its disk and to dissect its winds. The ultraviolet resonance = spectral lines indicate that this type of cataclysmic variable suffers a = high mass loss rate in the form of high velocity winds. The observations = were completed as scheduled, and no problems were reported. 1.7 Completed Two Sets of WF/PC-2 7203 (Cosmology with the Medium Deep = Survey - Short Orbit Cases) The WF/PC-2 was used to take short (about 20 minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems = were reported. 1.8 Completed WF/PC-2 6906 (WF/PC-2 Cycle 6 Calibration: Internal = Monitor-Visual Flats) The WF/PC-2 was used to make internal calibration visual flat = images as part of the Cycle 6 internal monitoring of the instrument's = operations. This proposal defines the standard series of internal flat = fields, which will be executed twice per month. The observations were = completed as planned, and no anomalies were reported. 1.9 Completed GHRS/2 6010 (Lyman Alpha Observations of High Radial = Velocity Stars) The High Resolution Spectrograph (Side Two) was used to make = ultraviolet spectrographic observations of the star HD6755 which is known = to have a high radial velocity in order to study its Lyman-Alpha line = complex. A wavelength calibration image was also taken. The = observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 6002 (Spectroscopy of the Rapidly Rotating = Nuclear Disk in the Virgo S0 NGC-4526) ' The WF/PC-2 was used to make spectrographic observations of the = rapidly rotating nuclear disk in S0 class galaxy NGC-4526 which is = located in the Virgo Cluster. The nucleus of this galaxy was imaged in = Cycle 4, and new images with different filters were used to make color = corrections and allow the subtraction of the stellar continuum. The = observations were completed as planned, and no problems were reported. 1.11 Completed Three Sets of WF/PC-2 6609 (The WF/PC-2 B-Band Parallel = Survey: A Systematic and Synoptic Study ) ' The WF/PC-2 was used to make B-Band survey observations in = parallel with the prime science observations. A total of 216 exposures = will be taken over a period of time to make a statistical study of the for= mation and evolution of the galaxies. The observations were executed as = scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTULIR09. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.13 Completed FOS/RD 6548 (The Cool Interstellar Medium in Hot = Cluster of Galaxies) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet resonance absorption line observations of quasar QSO0020+28 = in order to study the cool interstellar medium within a cluster of = galaxies which lie within the line of sight to the quasar. The QSO is = being used a probe of the galactic cluster. The observations were = executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 7202 (The WF/PC-2 B Band Parallel Survey: A = Systematic and Synoptic Study of Galaxy Formation and Evolution - Short = Orbit Case) The WF/PC-2 was used to make observations as part of a B-Band = Survey to make synoptic study of galaxy formation and evolution. These = observations were taken in parallel with the prime science observations. = They were completed as planned, and no anomalies were reported. 1.15 Completed GHRS/1 6463 (The Origin of the Peculiar Hybrid PG 1159 = Stars) The High Resolution Spectrograph (Side One) was used to make = observations of NGC-70 and an Abell Cluster in order to study the origin = of the peculiar hybrid STAR-2324 Class of stars. The data will be used = to obtain an improved value of the effective temperature, and the spectra = will be compared to the latest model atmospheres to learn more about the = evolution of PG 1159 stars. The observations were completed as = scheduled, and no problems were noted. 1.16 Completed FOS/RD 6577 (An Expanded Survey for Study of the = Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red Side) was used to make = observations of the QSO PKS1354+25 in order to make a survey of the = Lyman-Alpha line and the 68 Mg II absorption line system. The = observations were completed as scheduled, and no problems were reported. 1.17 Completed FOS/BL 6799 (The BAL Region Covering Factor in a Sample = of IRAS-Selected QSOs) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of QSO IR20450-295 in order to determine the = criteria by which the presence of BALs in the spectra may be predicted. = The observations were completed as planned, and no problems were = reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 19 Successful : 19 Scheduled Reacquisitions : 33 Successful : 33 2.2 FHST Updates: Scheduled : 29 Successful : 29 2.3 Operations Notes: There were no significant problems encountered over the weekend. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH